# Dispensing with optimal control: a new approach for the pricing and management of share buyback contracts

### v1

@@ Introduction
Assume that the price process $(S_{t})_{t}$ of the stock is given by
@@
The dynamic programming principle tells us that
@@
More precisely, the idea consists first in choosing a family
@@
For the first approach – hereafter denoted by PDE, we use an implicit Euler scheme for Equation (4) on $[0,T]\times\left[0.8,1.2\right]$ with Neumann conditions at the boundaries.
For the second approach – hereafter denoted by LS, we approximate the conditional expectations with polynomials of order 2 in the $y$ variable.
For the third approach – hereafter denoted by OHS, the family ${\cal F}$ is given by
@@
We then play the strategies obtained with the different methods using a Monte-Carlo simulation with $10,000$ trajectories. The average payoff and standard deviation for each method is reported in Table 1. The distribution of the values obtained for each method are plotted in Figures 1, 2, and 3, respectively. It is remarkable that the three strategies yield very close results both in terms of expected value and standard deviation. The distributions under $\mathbb{Q}$ of the resulting payoffs are spread below $1$ and have a spike above $1$. In all cases, the variability of the payoffs can be hedged away using standard mathematical finance techniques.
@@
Unlike in the previous section, we consider here a discrete-time model with daily decision making: on day $n$, the trader chooses the quantity $v_{n}\in[v_{min},v_{max}]$ to be bought, with $0\leq v_{min}<v_{max}$ specified in the contract.
@@
To incentivize the trader to spend at least the minimum required notional, [^7] At the stopping time, we buy the required number of shares to reach the minimum notional if it has not been reached yet. we slightly modify the payoff into
@@
We compare the different strategies, and report the results in the table below. Notice first that allowing the trader to buy more shares always results in an increased PnL. Again, our six strategies clearly beat the benchmarks.
@@
Optuna-$\alpha,a$ 2.84 142.19 94.25
Optuna-$\alpha,\beta,\gamma,a$ 2.88 143.75 98.39
Optuna-$\alpha,a,a_{1}$ 2.87 143.48 102.14
Optuna-$\alpha,a,a_{1},a_{2}$ 2.87 143.62 99.09
Optuna-$\alpha,a,a_{1},b$ 2.89 144.55 85.69
Optuna-$\alpha,a,a_{1},b,b_{1},c$ 2.90 145.20 80.81
@@
Linear policy 1.36 68.21 83.49
Minmaxtarget policy 2.19 109.59 99.48
@@
Optuna-$\alpha,a$ 2.79 139.30 92.69
Optuna-$\alpha,\beta,\gamma,a$ 2.81 140.52 95.84
Optuna-$\alpha,a,a_{1}$ 2.80 140.11 98.13
Optuna-$\alpha,a,a_{1},a_{2}$ 2.81 140.27 96.84
Optuna-$\alpha,a,a_{1},b$ 2.83 141.61 88.10
Optuna-$\alpha,a,a_{1},b,b_{1},c$ 2.86 143.21 77.23
@@
In conclusion, our research marks a significant shift from traditional approaches that relied on stochastic optimal control tools for pricing and managing share buyback contracts, which were originally inspired by optimal execution methods. By refining and optimizing heuristic strategies, we successfully address the limitations posed by high-dimensional state spaces and the complex selection of risk parameters. Our method simplifies the issue to pricing and hedging a Bermudan payoff, making it tractable with conventional financial tools. This approach not only facilitates a practical bridge between theory and application in financial markets but also separates the strategies for delivering shares to the company from the hedging of the payoff, potentially elucidating execution patterns observed in the market.
@@
This research has been conducted with the support of the Research Initiative “Modélisation des marchés actions, obligations et dérivés” financed by HSBC France under the aegis of the Europlace Institute of Finance. The views expressed are those of the authors and do not necessarily reflect the views or the practices at HSBC.

### v2

@@ Introduction
Assuming zero interest rates for the sake of simplicity, we consider that the price process $(S_{t})_{t}$ of the stock is given by
@@
The dynamic programming principle associated with the discretization of the problem over the above subdivision is written as follows:
@@
More precisely, the idea consists first in choosing a family of functions
@@
For the first approach, hereafter denoted by PDE, we use an implicit Euler scheme for Equation (4) on $[0,T]\times\left[0.8,1.2\right]$ with Neumann conditions at the boundaries.
For the second approach, hereafter denoted by LS, we approximate the conditional expectations with polynomials of degree $2$ in the $y$ variable.
For the third approach, hereafter denoted by OHS, the family ${\cal F}$ is given by
@@
We then test the strategies obtained with the different methods using a Monte-Carlo simulation with $10,000$ trajectories. The average payoff and standard deviation for each method is reported in Table 1. The distributions of the values obtained for the different methods are plotted in Figures 1, 2 and 3, respectively. It is remarkable that the three strategies yield very close results both in terms of expected value and standard deviation. The distributions under $\mathbb{Q}$ of the resulting payoffs are spread below $1$ and have a spike above $1$. In all cases, the variability of the payoffs can be hedged away using standard mathematical finance techniques.
@@
We consider here a discrete-time model with daily decision making: on day $n$, the trader chooses the quantity $v_{n}\in[v_{min},v_{max}]$ to be bought, with $0\leq v_{min}<v_{max}$ specified in the contract.
@@
To incentivize the trader to spend at least the minimum required notional, [^7] At the stopping time, we buy the required number of shares to reach the minimum notional if it has not been reached. we slightly modify the payoff into
@@
We compare the different strategies, and report the results in the table below. Notice first that allowing the trader to buy more shares always results in an increased PnL. Again, our six strategies clearly beat the benchmarks. We also see in this case with floating notional that increasing the complexity of the strategy by adding a dependence of the execution strategy on the $\frac{A}{S}$ ratio significantly enhances the average PnL.
@@
Optuna-$\alpha_{,}a$ 2.42 121.06 79.34
Optuna-$\alpha,\beta,\gamma,a$ 2.45 122.52 81.04
Optuna-$\alpha,a,a_{1}$ 2.44 122.12 81.03
Optuna-$\alpha,a,a_{1},a_{2}$ 2.45 122.31 80.54
Optuna-$\alpha,a,a_{1},b$ 2.75 137.68 91.71
Optuna-$\alpha,a,a_{1},b,b_{1},c$ 2.80 140.13 77.02
@@
Linear policy 0.24 11.75 74.57
Minmaxtarget policy 2.03 101.57 118.23
@@
Optuna-$\alpha,a$ 2.18 108.76 89.44
Optuna-$\alpha,\beta,\gamma,a$ 2.19 109.68 88.05
Optuna-$\alpha,a,a_{1}$ 2.20 109.83 85.85
Optuna-$\alpha,a,a_{1},a_{2}$ 2.20 109.86 89.09
Optuna-$\alpha,a,a_{1},b$ 2.75 137.42 98.32
Optuna-$\alpha,a,a_{1},b,b_{1},c$ 2.79 139.59 78.22
Again, our six strategies clearly beat the benchmarks. We also see, as above, that increasing the complexity of the strategy by adding a dependence of the execution strategy on the $\frac{A}{S}$ ratio significantly enhances the average PnL.
@@
In conclusion, our research marks a significant shift from traditional approaches that relied on stochastic optimal control tools for pricing and managing share buyback contracts, which were originally inspired by optimal execution methods. By refining and optimizing heuristic strategies, we successfully address the limitations posed by high-dimensional state spaces and the complex selection of risk parameters. Our method simplifies the issue to pricing and hedging a Bermudan payoff, making it tractable with conventional financial tools. This approach not only builds a bridge between theory and application in financial markets but also separates the strategies for delivering shares to the company from the hedging of the payoff, potentially elucidating execution patterns observed in the market.
@@
This research has been conducted with the support of the Research Initiative “Modélisation des marchés actions, obligations et dérivés” financed by HSBC France under the aegis of the Europlace Institute of Finance. Manuel Abellan-Lopez (HSBC), Jérôme Lemue (HSBC) and Stéphane Quentin (HSBC) also deserve special thanks for the discussions we had with them. The views expressed are those of the authors and do not necessarily reflect the views or the practices at HSBC.